OpenStudy (anonymous):

FAN AND MEDAL!!! Given f(x) = X+7/5, solve for f-1(3).

OpenStudy (anonymous):

ii. Rewrite -3x + 4y = 11 as a function of x. You must solve the given equation for y, since y = f(x). then (3x + 11)/4 = y gives f(x) = (3x + 11)/4. iii. Graph the function using the intercepts : y-intercept (x= 0) is y = -2(0) -5 = -5. So one point on the line is (0, -5). The x-intercept (y=0) is 0 = -2x - 5 => 2x = 5 => x = 5/2. iv. If f(x) = 4x - 3, find f(-1). Use the same technique as (i) above. Thus f(-1) = 4(-1) - 3 = -4-3 = -7.

OpenStudy (anonymous):

I think? but i think im rite:)

OpenStudy (anonymous):

SO the answer is -7?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

but thats not one of my options...

OpenStudy (anonymous):

well? im sorry than :I

OpenStudy (johnweldon1993):

Wait..are we solving "Given f(x) = X+7/5, solve for f-1(3)."

OpenStudy (solomonzelman):

that's what she said... the question was stated well enough !!

OpenStudy (johnweldon1993):

@SolomonZelman Correct but looking at ixerroks first post above, it looks like a completely different question! ..regardless

OpenStudy (solomonzelman):

bubbles are definitely cool, lol !

OpenStudy (solomonzelman):

Well, he is JUST WRONG.

OpenStudy (solomonzelman):

or SHE , who knows ;) ?

OpenStudy (anonymous):

lol yeah cx

OpenStudy (johnweldon1993):

\[\large y = x + \frac{7}{5}\] rewrite f^-1(x) means you find the inverse of this function...you do this by switching the x and 'y' variables and then solving again for 'y' \[\large x = y + \frac{7}{5}\] To solve for 'y' again...subtract 7/5 from both sides of the equation \[\large y = x - \frac{7}{5}\] Now we focus on the fact it asks for f^-1(3) ...we now plug in 3 for every 'x' we see in our equation... \[\large y = 3 - \frac{7}{5}\] \[\large y = \frac{15 - 7}{5}\] \[\large y = \frac{8}{5}\]

OpenStudy (anonymous):

Sorry then look like i got the wrong question (oops)

OpenStudy (anonymous):

its \[(x+7) / 5\]

OpenStudy (anonymous):

Oh. Sorry about that @bubbles-are-cool.

OpenStudy (anonymous):

its fine o:

OpenStudy (anonymous):

Thanks :) @bubbles-are-cool. I had just got the question mixed up :$

OpenStudy (johnweldon1993):

Ahh..changes things lol \[\large y = \frac{x + 7}{5}\] Again switch \[\large x = \frac{y + 7}{5}\] Multiply both sides by 5 \[\large 5x = y + 7\] subtract 7 from both sides \[\large y = 5x - 7\] NOW plug in x = 3 \[\large y = 5(3) - 7\] \[\large y = 15 - 7\] \[\large y = 8\]

OpenStudy (anonymous):

lol thanks so much!

OpenStudy (johnweldon1993):

No problem :)

OpenStudy (anonymous):

hey good job @johnweldon1993